What is the value of r in 2/3r = 6?
Answer by rfadrogane(214)

What is the value of r in 2/3r = 6?
Sol'n:
2/3r = 6
r = 6(3/2)
r = 9 ---answer
^_^
try our new tutorial site soon
mathgeas.com, the playground of great minds